OpenStudy (anonymous):

need help does anybody here is intersted? i appreciate a lot :) so i have this problem im struggling Six times Jason's collection of books and one-third of Nathan's collection add up to 134 books. One-third of Jason's collection and Nathan's entire collection add up to 31 books. what is the # of book for Jason Collection? and the # of books for Nathan Collection?

OpenStudy (perl):

ok

OpenStudy (bioepic):

Me 2.

OpenStudy (bioepic):

\[6 * ( j - ( 1/3 ) n ) = 134\] \[(1/3)j+n=31\] These are 2 equations u need to solve.

OpenStudy (anonymous):

hmmm ok......thank you

OpenStudy (bioepic):

No problem! Let me know what u get so I can check ur work! :) XD

OpenStudy (anonymous):

@perl

OpenStudy (perl):

first statement tells us 6*J + 1/3*N = 134

OpenStudy (perl):

then 1/3J + N = 31

OpenStudy (anonymous):

so it dont matter with which equation i start?...

OpenStudy (anonymous):

uhh i tried but nothing seem t make sence..so how do you exacly solve this equations?

OpenStudy (bioepic):

Do you know how to solve a system of equations?

OpenStudy (bioepic):

Since there are two variables, normal methods won't work.

OpenStudy (anonymous):

can't remeber ..do you subtitude or something?

OpenStudy (anonymous):

so tha would be j=21 and n=24?

OpenStudy (bioepic):

Yep! Nice work!

OpenStudy (anonymous):

uhh thank God u guys are king Thanks u so much..i see u soon jaja :)

OpenStudy (bioepic):

See ya!